Nick is given $50 to spend on a vacation. He decides to spend $5 a day. The amount Nick has left and the number of days are related. How many days Nick have after $25?


1
Expert's answer
2022-03-23T02:07:54-0400

Nick is given $50 to spend on a vacation. He decides to spend $5 a day. The relation of the amount of money M Nick has left and the number of days d would look like:

M = 50 - 5d (1)

If Nick has spent $25, then he is left with:

50 - 25 = 25

We can take relation 1 and knowing that amount of money M left equals to 25 we can find number of days left:

50 - 5d = 25

5d = 25

d = 5

So answer is after $25 Nick have 5 days left.